OpenStudy (anonymous):

Find the sum of the given arithmetic series. 24 + 48 + 72 + 96 + ... + 480

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ a_n=a_1+d(n-1) }\) and in this case, \(\normalsize\color{blue}{ 480=24+24(n-1) }\) \(\normalsize\color{blue}{ 480=24(1)+24(n-1) }\) \(\normalsize\color{blue}{ 480=24(n-1+1) }\) \(\normalsize\color{blue}{ 480=24(n) }\) \(\normalsize\color{blue}{ n= 20 }\)

OpenStudy (solomonzelman):

So, we found that there are 20 terms.

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ S_n=\frac{1}{2}~(a_1+a_n)\times n }\)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ S_{20}=\frac{1}{2}~(24+480)\times 20 }\) \(\LARGE\color{blue}{ S_{20}=(24+480)\times 10 }\) \(\LARGE\color{blue}{ S_{20}=(504)\times 10 }\) \(\LARGE\color{blue}{ S_{20}=5040 }\)

OpenStudy (solomonzelman):

So the answer is 5040.

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

find the sum of the given arithmetic series. (: 1.30+23+16+9+...(-30) 2. 26+19+12+5+...+(-37) 3.27+20+13+6+...+(-36)

OpenStudy (anonymous):

can you help me on these also? :)